Question from chemical and ionic equilibrium:
 
A vessel at 1000K has CO2 at 0.5 atm pressure. Some of the CO2 is  converted into CO on the addition of graphite. If the total pressure at equilibrium is 0.8atm, find the value of Kp
 
Please Help............

Grade:12

1 Answers

Arun
6 years ago
CO2 +C ------2COinitial presasure=.5atmfinal pressure on addition of graphiteCO2=.5-XCO=2Xtotal pressure at equilibrium=.5+xbut the total pressure at equilibrium= .8atm,5+x=.8atmx=.3atmthe value of reaction constant,K=[CO]^2/[CO2]=.6^2/.2=1.8  RegardsArun (askIITians forum expert)